tayle:

Dimitri is paid 8 percent commission on his total sales every pay period. Last pay period, he sold $6,300 worth of items. After a state tax of 6 percent is deducted from his commission, how much is left to the nearest dollar?

Imagine:

Dimitri is left with $474. Dimitri is paid 8% commission on his total sales every pay period. He sold $6300 worth of items so commission on the sold goods will be = 8% of $6300 = 6300× = 6300×0.08 = $504 A state tax of 6% has been deducted from the commission so we have to find out the amount left. Left amount = 504 - 6% of 504 = 504 - (504×.06) = 504 - 30.24 = $473.76 ≈ $474 Therefore, Dimitri is left with $474.

Hoodmemes:

Ok well this is simple U wanna make 8% a decimal \[8\div 100= 0.08\] \[0.08\times 6300=504\] Now the deduction \[6\div 100=0.06\] \[0.06\times 504=30.24\] Now subtractin \[504-30.24=473.76\] Final answer (743)